OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (-6, 4) and (2, 0). y - 4 = -one half(x + 6) y - 4 = 2(x + 6) y + 6 = -one half(x - 4) y + 6 = 2(x - 4)

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ -6}}\quad ,&{\color{blue}{ 4}})\quad &({\color{red}{ 2}}\quad ,&{\color{blue}{ 0}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}} \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (anonymous):

0-4=m(2-6)

OpenStudy (jdoe0001):

hmmm what did you get for "m"? or the slope

OpenStudy (anonymous):

never mind i found the answer